I would really appreciate your help.
how do I simplify (3 + i) ^2?
and how do I simplify square rt -81 * square rt -36?

Answer by sudhanshu_kmr(1152)


(3 + i) ^2 = 3^2 + 2*3*i + i^2
= 9 + 6i -1
= 8 + 6i

square rt -81 * square rt -36 = 9i * 6i
= 54 *i^2
= -54


Everyone is welcome to get online help: sudhanshu.cochin@gmail.com